Although the Pacific Ocean, Baltic Sea, and Arctic Sea represent three bodies of water, when added together, their sum yields seven Cs. Using the nine digits {0,2,3,4,5,6,7,8,9} substitute a different digit for each letter to make a true sum for PACIFIC + BALTIC + ARCTIC = CCCCCCC. Notice that 1 is not a possible digit for this problem. Your task is to provide the value of the word PACIFIC.
Answer by scott8148(5896)

starting with the ones digit ___ C+C+C=?C ___ C=5 is the only number that works, so a 1 is carried

tens digit ___ I+I+I+1=?5 ___ I=8 is the only number that works, so a 2 is carried

hundreds digit ___ T+T+F+2=?5 ___ 2 MUST be carried because of L value in thousands
___ so sum is 25, T=7 and F=9

thousands digit ___ 8+L+5+2=?5 ___ sum is 15 so L=0 ___ 1 is carried

ten-thousands digit ___ 5+A+R+1=?5 ___ sum is 15 ___ 1 is carried ___ A+R=9

hundred-thousands digit ___ A+B+A+1=?5 ___ sum is 15 ___ A=6 and B=2 ___ R=3

millions digit ___ P+1=5 ___ P=4